OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) x + 5y = 3 2x + 9y = 4

OpenStudy (anonymous):

this time solve the first equation for \(x\) via \(x=3-5y\) then substitute \(3-5y\) for \(x\) in the second equation to get \[2(3-5y)+9y=4\] or \[6-10y+9y=4\]\[6-y=4\]\[-y=-2\]\[y=2\]

OpenStudy (anonymous):

then substitute \(2\) for \(y\) in either equation to find \(x\)